OpenStudy (anonymous):

Write the standard equation of each Ellipse. Find the center, vertices, co-vertices and foci. 1. 25x2 + 9y2 = 225 2. 49x2 + y2 = 49 3. x2 - 2x + 9y2 - 8 = 0 4. x2 + 4y2 - 18x -8y + 81 = 0

OpenStudy (anonymous):

Familiar with 'completing-the-square?'

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok, we can review that. There is an equivalent way to do the same thing that might be easier, so I'll show you that too. You'll only need that for #3 and #4; the first two can be solved immediately.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

The standard form for an ellipse looks like this. \[\large \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The center is at (h,k) and a and b are the semi-axes. Look at #1 and see what you have to do to put it into standard form.

OpenStudy (anonymous):

. 25x2 + 9y2 = 225

OpenStudy (anonymous):

It needs to be set equal to 1, so what's the easy step that will take care of that?

OpenStudy (anonymous):

So that's all you have to do for #1 and #2; let me know when you are ready to move on to the last two.